OpenStudy (anonymous):

A student studying for a vocabulary test knows the meanings of 10 words from a list of 20 words. If the test contains 10 words from the study list, what is the probability that at least 8 of the words on the test are words that the student knows? (Round your answer to three decimal places.)

OpenStudy (kropot72):

The probability is found from: \[P(8knownWords)=\frac{\left(\begin{matrix}10 \\ 8\end{matrix}\right)\left(\begin{matrix}10 \\ 2\end{matrix}\right)}{\left(\begin{matrix}20 \\ 10\end{matrix}\right)}\] Do you know how to calculate this